# On the use of far-infrared and sub-millimeter line ratios to constrain the physical properties of gas in high-z galaxies

### v1

@@ 3.2 O352O388
The [OIII] m / [OIII] m (O352O388) ratio has been found to be an indicator for the gas density (Dinerstein et al., 1985; Rubin et al., 1994; Brauher et al., 2008; Nagao et al., 2011, 2013; Pereira-Santaella et al., 2017; Rigopoulou et al., 2018; Killi et al., 2023; Nakazato et al., 2023). Indeed, in contrast to the optical [OIII] lines (e.g. 4363, 4959, 5007), the FIR [OIII] line ratio is less dependent on the electron temperature and more sensitive to the electron density (Dinerstein et al., 1985).
The total O352O388 of Ponos is 0.36. Recent ALMA observations by Killi et al. (2023) found a ratio of 0.7 for a galaxy at z = 7, which is a factor of two higher than Ponos. Simulation work by Nakazato et al. (2023) found their simulated galaxies at z = 7 to have ratios , ranging from about 0.8 - 1.0. Our result, as well as the findings by Killi et al. (2023) and Nakazato et al. (2023) are lower than local galaxies (Brauher et al., 2008), and local planetary nebulae (Dinerstein et al., 1985). As shown in the second panel of Fig. 2, there are spatial variations in the map, with generally lower ratios in the inner regions of the disc and merger, which radially increase towards the outskirts. In the second row of Fig. 3, the strongest dependency can be seen on the gas density, where the scatter is also very small, which is in agreement with previous findings (Nagao et al., 2011; Pereira-Santaella et al., 2017). This is because the two [OIII] lines have slightly different critical densities ( cm-3 for [OIII]52 and cm-3 for [OIII]88 respectively) and thus their ratio is sensitive to the gas density. Higher densities lead to higher ratios, with an increase of four orders of magnitude from n = cm-3 to n = cm-3. The dependency on temperature is weak and degenerate, as higher O352O388 ratio values are associated with both low and high temperatures. All galaxy components show a completely flat trend of O352O388 as a function of gas metallicity, which is expected as both lines trace the same element. There is a dependency on the radiation field, where we see a rise in ratio towards higher radiation fields. The outflowing and inflowing component show a significant difference in ratio as a function of the velocity, where the ratio of the outflowing gas decreases by over an order of magnitudes. This indicates that the [OIII]88 line is a stronger tracer of outflows than the [OIII]52 emission line.
@@ 3.3 N2N3
The [NII] m / [NIII] m (N2N3) ratio can be used to study the ionisation parameter of the gas (Nagao et al., 2011), and to measure the effective temperature of stars ionising the gas (Rubin et al., 1994; Brauher et al., 2008). Singularly ionised nitrogen (N+) has an ionisation potential of 14.53 eV, and a critical density of cm-3, and thus [NII] traces lower temperature areas in ionised HII regions than doubly ionised nitrogen (N++), which has an ionisation potential of 29.60 eV.
From the third panel of Fig. 2, it can be seen that, in contrast to the previously discussed ratios, the N2N3 ratio is almost uniform across the source, with an enhancement in small areas of lower temperature ionised gas (104 K ¡ T ¡ 104.5 K), where [NII] dominates. In the third row of Fig. 3, we see that the N2N3 ratio is degenerate with gas density. The most significant dependency of this ratio is on the gas temperature, where we see higher temperatures leading to lower ratios. The decrease of N2N3 as a function of temperature is steeper for disc and merger components than for the CGM, because of the low radiation field strength of the CGM. As metals in gas promote the cooling, we investigated if the ability of nitrogen to trace the temperature of the ionised gas could be affected by a change in metallicity. We found no significant differences in the dependency of the N2N3 ratio on the temperature at different metallicities.
The N2N3 ratio does not depend on the metallicity, until we reach super-solar values, where the ratio rapidly drops, which can once again be connected to hot gas that has been enriched by SNe, and experiences a strong radiation field, which is the same trend we see in the corresponding increase of O3C2. There is also a dependency on the radiation field, with higher ratios for lower radiation fields, which is caused by the lower ionisation energy of [NII] compared to [NIII]. We do not detect a significant difference in N2N3 ratio between inflows and outflows. Due to the weak dependencies on density and metallicity, we find that N2N3 can be used as a diagnostic tool to constrain the ionised gas temperature. We also find a dependency on the radiation field, and thus the ionisation of the gas, which is in agreement with Nagao et al. (2011), who found that N2N3 traces the ionisation parameter.
@@ 3.4 C2N2
Like the O3C2 ratio, the [CII] m / [NII] m (C2N2) ratio uses two emission lines that originate from different phases of the ISM. The [CII] and [NII] lines have ionisation energies of 11.26 eV and 14.53 eV respectively, which are both close to the ionisation energy of hydrogen (13.6 eV). Previous studies have tested the [CII] m / [NII] m (or [NII]205/[CII]) ratio, and proposed it to be a tracer of metallicity (Nagao et al., 2013), and as a tracer to constrain the fraction of [CII] emission that is associated with ionised gas (Decarli et al., 2014) and to scale with the IR-luminosity (Cunningham et al., 2020). The ratio of [CII] and [NII]122 has not yet been well explored.
The spatially resolved map of this ratio (bottom-left panel of Fig. 2) shows higher values in dense structures of the disc. The plots reported in Fig. 3 (fourth row) shows evidence for a phase transition of C2N2 as a function of density and temperature. In the low density regime, the ratio has a slight tendency to decrease with increasing density until it reaches a turning point at n = 5 cm-3, above which the ratio rapidly increases by two orders of magnitude with density. [CII] has a wide range of critical densities for for collisions with electrons, atoms or molecules (5 cm-3 up to 103 cm-3), and the C2N2 ratio starts to increase when the minimum critical density is reached. The same behaviour can be seen in the temperature, where the ratio quickly drops off by two orders of magnitude until it reaches a temperature slightly above 104 K, after which it remains roughly constant with temperature. Therefore, denser and colder regions of the ISM are more sensitive to the C2N2 ratio than diffuse gas, which can also be seen in the map from Fig. 2. The turning point in both temperature and density shows evidence for a phase transition form ionised HII regions, to neutral PDR regions. Therefore we find that the C2N2 ratio could be used to infer the gas density and temperature of PDR regions, as it seems to be most sensitive in this regime. On the other hand, the C2N2 ratio is insensitive to the gas phase metallicity. This insensitivity would not significantly change if we would directly trace the abundances of elements, as the C/N ratio has been found to be approximately constant over a wide range of metallicities (Garnett et al., 1999; Berg et al., 2016; Pereira-Santaella et al., 2017). The ratio is also only slightly sensitive to the radiation field, where a large scatter within the disc component can be seen. In the merger and the disc the ratio increases towards lower radiation fields, which corresponds to shielded, denser gas. There is a general weak trend for lower ratios at higher radiation fields. We see no differences in C2N2 ratio between the outflowing and the inflowing gas.
@@ 3.5 O3N3
The [OIII] m / [NIII] m (O3N3) ratio is a candidate for tracing the gas metallicity, as it is a good proxy for the N/O abundance ratio (Nagao et al., 2011; Pereira-Santaella et al., 2017). Both lines have similar ionisation energies, and thus trace the same gas in HII regions, which is important for metallicity calibrations (Méndez-Delgado et al., 2023). To account for the dependence of the [OIII] lines on the electron density of the gas, the O3N3 ratio often combines the [OIII]88 and the [OIII]52 line (Pereira-Santaella et al., 2017). In our analysis we found that including the [OIII]52 line does not change our results, and thus we focus on the ratio including only the [OIII]88 line (which is easier to observe from ground). As discussed above for the O3C2 ratio, the assumption of constant abundance ratios in our modelling limits our capability of exploring true metallicity dependencies.
The second panel of the second row of Fig. 2 shows the O3N3 ratio map. As both lines originate from similar phases of the gas, the spatial variation of the ratio is small. The O3N3 ratio is higher in the inner regions of the disc, with small areas with highly increased ratios within the ISM, which correspond to hot and diffuse regions like HII regions and SNe sites, where [OIII] dominates. In the bottom row of Fig. 3 we can see a strong dependency of the ratio on the gas density, whereby the ratio decreases by over two orders of magnitude within the explored density range. The trend with the temperatures shows a similar behaviour, where gas with lower temperatures also has a lower ratio, and flattening towards higher temperatures. The radiation field seems to be the strongest dependency we find, where high radiation fields lead to a low ratio, rising by two orders of magnitude. Interestingly, we see a significant difference in ratio between outflowing and inflowing gas, whereby for the former, the O3N2 ratio increases by an order of magnitude as a function of velocity.
@@ 4 Conclusions
In this work we investigated FIR emission line ratios in the context of a cosmological zoom-in simulation of a high-redshift galaxy, to asses how much these ratios can tell us about the galaxy’s ISM properties. Our main results can be summarised as follows:
We discovered that C2N2 can be used as a tracer to infer the density and temperature of PDR regions;
O3O3 is confirmed to be a good tracer for the gas phase density, although we also find a dependency on the incident radiation field;
We find a significant difference of O3C2, O3O3 and O3N3 in in- and outflowing gas, suggesting that these ratios could be used to analyse high velocity components. [OIII]88 m is sensitive to high velocity outflowing gas;
N2N3 is found to be a good tracer of the temperature of ionised gas and confirmed to trace the radiation field in the ISM of galaxies;
@@ Acknowledgements.
This project has received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No 951815 (AtLAST). We thank the anonymous referee for providing insightful comments that helped us improve the analysis and interpretation of the results. The simulations were performed using the resources from the National Infrastructure for High Performance Computing and Data Storage in Norway, UNINETT Sigma2, allocated to Project NN9477K. We acknowledge usage of the Python programming language (Van Rossum & Drake Jr, 1995; Van Rossum & Drake, 2009), Matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011), and Pynbody (Pontzen et al., 2013).References
@@
- Fujimoto et al. (2024) Fujimoto, S., Ouchi, M., Nakajima, K., et al. 2024, ApJ, 964, 146
@@
Appendix A High metallicity gas
As stated in subsection 3.1, we investigate the reason behind the increase of the O3C2 ratio at super-solar metallicities in more detail. Figure 4 shows the phase diagram of the gas emitting in both [CII] and [OIII], with the metallicity on the x-axis and the temperature on the y-axis. As can be seen there is a wide range of temperatures at high metallicity, reaching above 105 K. This high temperatures gas also shows high emission line ratios. Additionally, there is a small region within the phase diagram with very high metallicities, that resides within a small temperature range of around 2104 K, which also displays very high ratios. This gas experiences a high FUV radiation field (log10(G0) 2 - 2.5) and drives the high metallicity ratio in Fig. 3.

### v2

@@ 3.2 O352O388
The [OIII] m / [OIII] m (O352O388) ratio has been found to be an indicator for the gas density (Dinerstein et al., 1985; Rubin et al., 1994; Brauher et al., 2008; Nagao et al., 2011, 2013; Pereira-Santaella et al., 2017; Rigopoulou et al., 2018; Killi et al., 2023; Nakazato et al., 2023). In contrast to the optical [OIII] lines (e.g. 4363, 4959, and 5007), the FIR [OIII] line ratio is indeed less dependent on the electron temperature and more sensitive to the electron density (Dinerstein et al., 1985).
The total O352O388 of Ponos is 0.36. Recent ALMA observations by Killi et al. (2023) found a ratio of 0.7 for a galaxy at z = 7, which is higher by a factor of two than Ponos. Simulation work by Nakazato et al. (2023) found their simulated galaxies at z = 7 to have ratios , ranging from about 0.8 - 1.0. Our result, as well as the findings by Killi et al. (2023) and Nakazato et al. (2023), is lower than results for local galaxies (Brauher et al., 2008) and local planetary nebulae (Dinerstein et al., 1985). As shown in the second panel of Fig. 2, there are spatial variations in the map, with generally lower ratios in the inner regions of the disc and merger. The ratios increase radially towards the outskirts. In the second row of Fig. 3, the strongest dependence can be seen on the gas density, where the scatter is also very small, which agrees with previous findings (Nagao et al., 2011; Pereira-Santaella et al., 2017). This is because the two [OIII] lines have slightly different critical densities ( cm-3 for [OIII]52 and cm-3 for [OIII]88), and thus, their ratio is sensitive to the gas density. Higher densities lead to higher ratios, with an increase of four orders of magnitude from n = cm-3 to n = cm-3. The dependence on temperature is weak and degenerate because higher O352O388 ratio values are associated with both low and high temperatures. All galaxy components show a completely flat trend of O352O388 as a function of gas metallicity, which is expected because the two lines trace the same element. There is a dependence on the radiation field, where we see a rise in the ratio towards higher radiation fields. The outflowing and inflowing component show a significant difference in ratio as a function of the velocity, where the ratio of the outflowing gas decreases by over an order of magnitudes. This indicates that the [OIII]88 line is a stronger tracer of outflows than the [OIII]52 emission line.
@@ 3.3 N2N3
The [NII] m / [NIII] m (N2N3) ratio can be used to study the ionisation parameter of the gas (Nagao et al., 2011) and to measure the effective temperature of stars that ionise the gas (Rubin et al., 1994; Brauher et al., 2008). Singly ionised nitrogen (N+) has an ionisation potential of 14.53 eV and a critical density of cm-3, and thus, [NII] traces areas with a lower temperature in ionised HII regions than doubly ionised nitrogen (N++), which has an ionisation potential of 29.60 eV.
The third panel of Fig. 2 shows that in contrast to the previously discussed ratios, the N2N3 ratio is almost uniform across the source, with an enhancement in small areas of lower-temperature ionised gas (104 K ¡ T ¡ 104.5 K), where [NII] dominates. In the third row of Fig. 3, the N2N3 ratio is degenerate with the gas density. The most significant dependence of this ratio is on the gas temperature, where higher temperatures lead to lower ratios. The decrease of N2N3 as a function of temperature is steeper for disc and merger components than for the CGM because of the radiation field strength of the CGM is low. As metals in gas promote the cooling, we investigated whether the ability of nitrogen to trace the temperature of the ionised gas could be affected by a change in metallicity. We found no significant differences in the dependence of the N2N3 ratio on the temperature at different metallicities.
The N2N3 ratio does not depend on the metallicity until supersolar values are reached, where the ratio rapidly drops. This can once again be connected to hot gas that has been enriched by SNe and experiences a strong radiation field. This is the same trend as in the corresponding increase of O3C2. There is also a dependence on the radiation field, with higher ratios for lower radiation fields. This is caused by the lower ionisation energy of [NII] compared to [NIII]. We do not detect a significant difference in N2N3 ratio between inflows and outflows. Due to the weak dependences on density and metallicity, we find that N2N3 can be used as a diagnostic tool for constraining the ionised gas temperature. We also find a dependence on the radiation field, and thus, on the ionisation of the gas, which agrees with Nagao et al. (2011), who found that N2N3 traces the ionisation parameter.
@@ 3.4 C2N2
Like the O3C2 ratio, the [CII] m / [NII] m (C2N2) ratio uses two emission lines that originate in different phases of the ISM. The [CII] and [NII] lines have ionisation energies of 11.26 eV and 14.53 eV, respectively, which are both close to the ionisation energy of hydrogen (13.6 eV). Previous studies have tested the [CII] m / [NII] m (or [NII]205/[CII]) ratio. It was proposed to be a tracer of metallicity (Nagao et al., 2013) and as a tracer for constraining the fraction of [CII] emission that is associated with ionised gas (Decarli et al., 2014). It was also suggested scale with the IR luminosity (Cunningham et al., 2020). The ratio of [CII] and [NII]122 has not been well explored so far.
The spatially resolved map of this ratio (bottom left panel of Fig. 2) shows higher values in dense structures of the disc. The plots reported in Fig. 3 (fourth row) show evidence for a phase transition of C2N2 as a function of density and temperature. In the low-density regime, the ratio has a slight tendency to decrease with increasing density until it reaches a turning point at n = 5 cm-3, above which the ratio rapidly increases by two orders of magnitude with density. [CII] has a wide range of critical densities for collisions with electrons, atoms, or molecules (5 cm-3 up to 103 cm-3), and the C2N2 ratio starts to increase when the minimum critical density is reached. The same behaviour can be seen in the temperature, where the ratio quickly drops off by two orders of magnitude until it reaches a temperature slightly above 104 K, after which it remains roughly constant with temperature. Therefore, denser and colder regions of the ISM are more sensitive to the C2N2 ratio than diffuse gas, which is also visible in Fig. 2. The turning point in both temperature and density shows evidence for a phase transition from ionised HII regions to neutral PDR regions. Therefore, we find that the C2N2 ratio could be used to infer the gas density and temperature of PDR regions, as it seems to be most sensitive in this regime. On the other hand, the C2N2 ratio is insensitive to the gas-phase metallicity. This insensitivity would not significantly change if we were able to directly trace the abundances of elements, as the C/N ratio has been found to be approximately constant over a wide range of metallicities (Garnett et al., 1999; Berg et al., 2016; Pereira-Santaella et al., 2017). The ratio is also only slightly sensitive to the radiation field, where a large scatter within the disc component can be seen. In the merger and the disc, the ratio increases towards lower radiation fields, which corresponds to shielded, denser gas. There is a general weak trend for lower ratios in higher radiation fields. The C2N2 ratio remains constant between the outflowing and inflowing gas.
@@ 3.5 O3N3
The [OIII] m / [NIII] m (O3N3) ratio is a candidate for tracing the gas metallicity because it is a good proxy for the N/O abundance ratio (Nagao et al., 2011; Pereira-Santaella et al., 2017). The two lines have similar ionisation energies and thus trace the same gas in HII regions, which is important for metallicity calibrations (Méndez-Delgado et al., 2023). To account for the dependence of the [OIII] lines on the electron density of the gas, the O3N3 ratio often combines the [OIII]88 and the [OIII]52 line (Pereira-Santaella et al., 2017). In our analysis, we found that including the [OIII]52 line does not change our results, and thus, we focused on the ratio that only included the [OIII]88 line (which is easier to observe from the ground). As discussed above for the O3C2 ratio, the assumption of constant abundance ratios in our modelling limits our capability of exploring the true metallicity dependences.
The second panel in the second row of Fig. 2 shows the O3N3 ratio map. The two lines originate from similar phases of the gas, and the spatial variation of the ratio is therefore small. The O3N3 ratio is higher in the inner regions of the disc. Small areas with highly increased ratios within the ISM correspond to hot and diffuse regions, such as HII regions and SNe sites, where [OIII] dominates. The bottom row of Fig. 3 shows a strong dependence of the ratio on the gas density, whereby the ratio decreases by over two orders of magnitude within the explored density range. The trend with temperature shows a similar behaviour, where gas with lower temperatures also has a lower ratio, and it flattens towards higher temperatures. The radiation field seems to be the strongest dependence we find, where high-radiation fields lead to a low ratio that rises by two orders of magnitude. Interestingly, the ratio of the outflowing and inflowing gas is different, whereby the O3N3 ratio increases by an order of magnitude as a function of velocity for the former.
@@ 4 Conclusions
We investigated FIR emission line ratios in the context of a cosmological zoom-in simulation of a high-redshift galaxy to asses the information provided by these ratios about the ISM properties of the galaxy. Our main results are listed below.
We discovered that C2N2 can be used as a tracer to infer the density and temperature of PDR regions.
O3O3 is confirmed to be a good tracer for the gas-phase density, although we also find a dependence on the incident radiation field.
We find a significant difference of O3C2, O3O3, and O3N3 in in- and outflowing gas, suggesting that these ratios could be used to analyse high-velocity components. [OIII]88 m is sensitive to high-velocity outflowing gas.
N2N3 is found to be a good tracer of the temperature of ionised gas and is confirmed to trace the radiation field in the ISM of galaxies.
@@ Acknowledgements.
This project has received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No 951815 (AtLAST). The simulations were performed using the resources from the National Infrastructure for High Performance Computing and Data Storage in Norway, UNINETT Sigma2, allocated to Project NN9477K. We thank the anonymous referee for providing insightful comments that helped us improve the analysis and interpretation of the results. We acknowledge usage of the Python programming language (Van Rossum & Drake Jr, 1995; Van Rossum & Drake, 2009), Matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011), and Pynbody (Pontzen et al., 2013).References
@@
- Fujimoto et al. (2022) Fujimoto, S., Ouchi, M., Nakajima, K., et al. 2022, arXiv e-prints, arXiv:2212.06863
@@
- Kumari et al. (2023) Kumari, N., Smit, R., Leitherer, C., et al. 2023, arXiv e-prints, arXiv:2307.00059
@@
Appendix A High metallicity gas in O3C2
We investigate the reason behind the increase of the O3C2 ratio at super-solar metallicities in more detail. Figure 4 shows the phase diagram of the gas emitting in both [CII] and [OIII], with the metallicity on the x-axis and the temperature on the y-axis. As can be seen there is a wide range of temperatures at high metallicity, reaching above 105 K. This high temperatures gas also shows high emission line ratios. Additionally, there is a small region within the phase diagram with very high metallicities, that resides within a small temperature range of around 2104 K, which also displays very high ratios. This gas experiences a high FUV radiation field (log10(G0) 2 - 2.5) and drives the high metallicity ratio in Fig. 3.